OpenStudy (anonymous):

The average of two numbers is WX. IF the first number is X what is the other number?

OpenStudy (sriram):

WX means W*X?

OpenStudy (bahrom7893):

(Y + X)/2 = WX Y + X = 2WX Y = 2WX-X Y = X(2W-1)